OpenStudy (haleyelizabeth2017):

What is the vertex form of the equation? Please just walk me through helping me :/ Thanks!

OpenStudy (haleyelizabeth2017):

\(y=-x^2+12x-4\)

OpenStudy (haleyelizabeth2017):

I know that I need to take out the -1: \(y=-1(x^2-12x+4)\)

OpenStudy (mertsj):

Factor -1 out of the first two terms on the right.

OpenStudy (haleyelizabeth2017):

huh?

OpenStudy (mertsj):

y=-1(x^2-12x + )+4

OpenStudy (haleyelizabeth2017):

how did you get the +4 out of the parenthesis?

OpenStudy (mertsj):

Make that +12x

OpenStudy (mertsj):

Would it be ok to group the terms like this: y=(-x^2+12x)-4

OpenStudy (haleyelizabeth2017):

I don't know :/

OpenStudy (mertsj):

It would. And having done so, factor -1 out of the first two terms: y=-1(x^2-12x)-4

OpenStudy (haleyelizabeth2017):

Okay

OpenStudy (mertsj):

Now complete the square inside the parentheses by adding 36: y=-1(x^2-12x+36)-4+36

OpenStudy (haleyelizabeth2017):

okay one second...I need to write this down and show the work on that part....one second :)

OpenStudy (mertsj):

Now factor and simplify: y=-1(x-6)^2+32 And that is vertex form

OpenStudy (haleyelizabeth2017):

okay thank you so much!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

Here,

OpenStudy (haleyelizabeth2017):

I figured it out....Thank you anyways! I have to explain a bunch of stuff so I have to write a ton of work so I prefer the longer way :D

OpenStudy (haleyelizabeth2017):

Can you help me with one more @Mirahalfian? I just got confused....I solved it but I just need to see if I did it right, so I need someone to walk me through the steps.....

OpenStudy (anonymous):

Ok sure 😊

TheSmartOne (thesmartone):

Sorry I wasn't online when you messaged me.

OpenStudy (haleyelizabeth2017):

it's okay :)